OpenStudy (anonymous):

How do you find two coterminal angles of an initial angle such as (pi/7)

OpenStudy (anonymous):

Coterminal angles are basically "the same". For example, pi/4 is the same as saying 9pi/4 because when you go all around, it brings you to the same spot. And so both of these angles are coterminal. Basically, you add or subtract 2pi (360 degrees), to find another angle that is coterminal with the given angle.

OpenStudy (anonymous):

So to find coterminal angles of pi/7, add/subtract 2pi:\[\bf \frac{ \pi }{ 7 }+2\pi=\frac{ 15\pi }{ 7 } \ or \ \frac{ \pi }{ 7 }-2\pi=-\frac{ 13\pi }{ 7 }\]

OpenStudy (anonymous):

Do you still need help?